OpenStudy (anonymous):

please help! cube route 24 + cube route 81 -cube route 54. i know i need to make the isdie of the route the same, but i don't know what to make it because theyre not all divisable by two

OpenStudy (anonymous):

Run through the first couple of perfect cubes: 2^3=8 3^3=27 This is probably as high as we will need. For each cube root we'll be looking to pull out one of these values because they are perfect cubes

OpenStudy (anonymous):

If you can factor out one of these perfect cubes, you can bring it outside the radical by cube rooting it. For example, look at the first piece, the cube root of 24: \[\sqrt[3]{24} = \sqrt[3]{8 \times 3} = 2\sqrt[3]{3}\]

OpenStudy (anonymous):

So the answer is 5 cube route 3 - 3 cube route 2?!

OpenStudy (anonymous):

Thats what I got, exactly

OpenStudy (anonymous):

okay thank you so much for your help!!

OpenStudy (anonymous):

No problem, glad I could help